Check the following equations for dimensional consistency where t is time (s), v is speed (m/s), a is acceleration (m/s2 ), and x is position (m): a) ? = ? 2 2? b) ? = 1 2 ?? c) ? = √ 2? ? .
Answer #1

answer) here we have dimensions

x=L

t=T

v=L/T

a=L/T2

so for a ) L=L/T*L/T2

L\neqL2/T3

so its dimensionally inconsistent

b) here we have

L=L/T2*T

L\neqL/T

so its dimensionally inconsistent

c) here we have

T=L*L/T2

T\neqL2/T2

so its dimensionally inconsistent